Limit n tends to infinity product r=3 to n(r^3-8)/(r^3+8)
 
 

Grade:12th pass

1 Answers

Aditya Gupta
5 years ago
(r^3-8)/(r^3+8)= (r-2)(r^2+2r+4)/(r+2)(r^2 – 2r+4)
=2/7*(n^2+3)((n+1)^2+3)/(n+2)(n+1)n(n-1)
so limit n tends to infinity product r=3 to n(r^3-8)/(r^3+8)= 2/7